a rectangle is given. perimeter is 108.the width is x and the length is 2x.calculate width and calculate the area
Answer by josgarithmetic(39614)

2(x+2x)=108
3x=54
x=27*2/3
x=18

x*2x=2x^2=2*18^2=648